# On Neural Networks as Infinite Tree-Structured Probabilistic Graphical Models

### v3

@@ 4 Implications and Extensions
We are not claiming that one should actually carry out the PGM construction used in the preceding section, since that PGM is infinite. Rather, its contribution is to let us understand precisely the approximation that SGD in a DNN is making; although a DNN itself can be understood as a BN or MN, SGD is not using that BN or MN but rather the infinite tree-structured one. While that PGM is infinite, is it built using the original as a template in a straightforward fashion, and hence is easy to understand. Beyond this contribution to comprehensibility and pedagogy, are there other applications?
@@
One application is an ability to use standard PGM algorithms such as Markov chain Monte Carlo (MCMC) to sample latent variables given observed values of input and output variables, such as for producing confidence intervals or understanding relationships among variables. One could already do so using Gibbs sampling in the BN or MN directly represented by the DNN itself (which we will call the “direct PGM”), but then one wouldn’t be using the BN or MN that SGD in the DNN actually used during training. For that, our result has shown that one instead needs to use Gibbs sampling in the infinite tree-structured PGM, which is impractical. Nevertheless, for any variable $V$ in the original DNN, on each iteration a Gibbs sampler takes infinitely many samples of $V$ given infinitely many samples of each of the members of $V$’s Markov blanket in the original DNN. By treating the variables of the original DNN as continuous, with their values approximating their sampled probabilities in the Gibbs sampler, we can instead apply Hamiltonian Monte Carlo or other MCMC methods for continuous variables in the much smaller DNN structure. We explore this approach empirically rather than theoretically in the next section. Another, related application of our result is that one could further fine-tune the trained DNN using other PGM algorithms, such as contrastive divergence. We also explore this use in the next section.
@@
One might object that most results in this paper use sigmoid activation functions. Nair and Hinton showed that rectified linear units (ReLU) can be thought of as a combination of infinitely many sigmoid units with varying biases (Nair and Hinton, 2010). Hence our result in the previous section can be extended to ReLU activations by the same argument. More generally, with any non-negative activation function that can yield values greater than one, while our BN argument no longer holds, the MN version of the argument can be extended. An MN already requires normalization to represent a probability distribution. While Batch Normalization and Layer Normalization typically are motivated procedurally, to keep nodes from “saturating,” and consequently to keep gradients from “exploding” or “vanishing,” as the names suggest, they also can be used to bring variables into the range $[0,1]$ and hence to being considered as probabilities. Consider an idealized variant of these that begins by normalizing all the values coming from a node $h$ of a neural network, over a given minibatch, to sum to $1.0$; the argument can be extended to a set of $h$ and all its siblings in a layer (or other portion of the network structure) assumed to share their properties. It is easily shown that if the parents of any node $h$ in the neural network provide to $h$ approximate probabilities that those parent variables are true in the distribution defined by the Markov network given the inputs, then $h$ in turn provides to its children an approximate probability that $h$ is true in the distribution defined by the Markov network given the inputs. Use of Batch or Layer Normalization is only approximate and hence adds an additional source of approximation to the result of the preceding section. Detailed consideration of other activation functions is left for further work; in the next section we return to the sigmoid case.
5 Alternative Training Algorithms: The Sigmoid Case
@@
To illustrate the potential utility of the infinite tree-structured PGM view of a DNN, in this section we pursue one of its potential implications in depth. We have already noted we can view forward propagation in an all-sigmoid DNN as exact inference in a tree-structured BN, such that the CPD of each hidden variable is a logistic regression. In other words, each hidden node is a Bernoulli random variable, with parameter $\lambda$ being a sigmoid activation (i.e. logistic function) applied to a linear function of the parent nodes. This view suggests alternative learning algorithms such as contrastive divergence (CD) that use sampling methods for high-dimensional binary random variables such as Gibbs sampling. Doing so has a natural advantage over SGD, which samples the values of the hidden variables using only the evidence at the input values. Instead, Gibbs sampling uses all the available evidence, both at input and output variables. MCMC has now advanced beyond Gibbs sampling with methods such as Hamiltonian Monte Carlo (HMC), but HMC will sample values in $\{0,1\}$ rather than in $[0,1]$.
To use HMC as proposed, we define hidden variables using the recently-developed continuous Bernoulli distribution (Loaiza-Ganem and Cunningham, 2019), where the single parameter $\lambda$ of the distribution is defined as in the Bernoulli case. Whereas the Bernoulli density is unnormalized when viewed (improperly) as a density over $(0,1)$, the continuous Bernoulli distribution is a proper density. Somewhat counter-intuitively, the expected value of this distribution is not equal to the parameter $\lambda$, which has important implications for inference. This option leads to learning algorithms that are able to take advantage of sampling methods effective for high-dimensional continuous random variables, including HMC.
With respect to this BN, whose variables correspond exactly with those of the DNN, we can see that our previous CD-1 algorithm, which is standard SGD for DNNs, samples settings of the input variables, computes expectations on all the remaining variables (latent and output variables), and adjusts the weights (CPDs) toward maximizing the probability conditional log likelihood, or minimizing cross-entropy. If instead of one gradient step we continued to convergence, the resulting algorithm almost would be the well-known Expectation Maximization (EM) algorithm for training any BN’s parameters from a data set with missing or hidden variables, given the BN structure. We say almost because this analysis reveals one important “error” or shortcoming in the SGD algorithm: in the E step where we compute the expected values of the latent variables, it entirely ignores the values of the output variables. In other words, we can view SGD as an approximation to EM in which evidence from output variables is ignored during the E step, and therefore gradients must be backpropagated across all layers to account for this evidence when updating the weights in the M step. This strategy is effective in later layers, which are closer to the output evidence, but highly ineffective in earlier layers. This limitation has been recognized as the vanishing gradient problem and addressed through ad-hoc initialization and pre-training strategies as well as Nesterov momentum. Nesterov momentum can be seen as “looking ahead” one step of SGD when computing the gradient, which partially incorporates evidence at output variables into the expectations at the hidden nodes.
More precisely and formally correcting this shortcoming is not easy: computing the correct expectation is NP-complete, and the most obvious algorithm always requires time exponential in the number of latent variables. In such situations in PGMs, researchers have replaced the E step with MCMC sampling (e.g., Gibbs) (Hinton et al., 2006; Song et al., 2016). However, running these MCMC chains to convergence is impractical, therefore it is common in practice to take a small number $k$ of steps in the chain between gradient steps, which again gives rise to the CD-1 or CD-$k$ algorithm. However, a different training algorithm for DNNs, motivated by the natural correspondence between DNNs and BNs described above – and which correctly accounts for evidence from the output variables through proper EM updates – converges to the correct answer in fewer training epochs compared to SGD.
The Continuous Bernoulli Bayes net (CBBN) is similar to the sigmoid BN (i.e., “stacked logistic regression"), but with logistic regression CPDs replaced by their continuous Bernoulli analogues. Equivalently, it is a feedforward, stochastic neural network in which hidden variables are continuous Bernoulli distributed. Consider a Bayesian network composed of input variables $\bm{x}=\bm{h}_{0}$, a sequence of layers of hidden variables $\bm{h}_{1},...,\bm{h}_{L}$, and output variables $\bm{y}$. Each pair of consecutive layers forms a bipartite subgraph of the network as a whole, and the variables $\bm{h}_{i}=(h_{i1},...,h_{iM_{i}})$ follow a multivariate continuous Bernoulli distribution with parameters $\bm{\lambda}_{i}=(\lambda_{i1},...,\lambda_{iM_{i}})$ that depend on variables in the previous layer $\bm{h}_{i-1}$ as follows:
$h_{ij}\sim\mathcal{CB}(\lambda_{ij})\text{, where }\bm{\lambda_{i}}=\sigma(\bm{W}_{i-1}\bm{h}_{i-1}+\bm{b}_{i-1}).$ (1)
$\sigma:\mathbb{R}\to(0,1)$ is a non-linearity – here the logistic function – that is applied element-wise, and $\bm{\theta}_{i}=(\bm{W}_{i},\bm{b}_{i})$ are parameters to be learned. For a complete setting of the variables $\{\bm{x},\bm{h},\bm{y}\}$, where $\bm{h}=\{\bm{h}_{1},...,\bm{h}_{L}\}$, and parameters $\bm{\theta}=\{\bm{\theta}_{i}\}_{i=0}^{L}$, the likelihood $p(\bm{y},\bm{h}|\bm{x};\bm{\theta})$ may be decomposed as:
$p(\bm{y},\bm{h}|\bm{x};\bm{\theta})=p(\bm{y}|\bm{h}_{L};\bm{\theta}_{L})\cdot\prod_{i=1}^{L}\prod_{j=1}^{M_{i}}p_{\mathcal{CB}}(h_{ij}|\lambda_{ij}(\bm{h}_{i-1};\bm{\theta}_{i-1})),$ (2)
where $p_{\mathcal{CB}}(\cdot|\cdot)$ denotes the continuous Bernoulli density, and a specific form for $p(\bm{y}|\bm{h_{L}},\bm{\theta_{L}})$ has been omitted to allow variability in the output variables. In our experiments, $\bm{y}$ is a Bernoulli or categorical random variable parameterized via the logistic or softmax function, respectively.
@@ 5.1 Learning via Contrastive Divergence with Hamiltonian Monte Carlo Sampling
Let $\bm{h}^{(0)},\bm{h}^{(1)},\bm{h}^{(2)},...$ denote a chain of MCMC samples of the complete setting of hidden variables in our CBBN. As previously noted, we allow hidden variables $h_{ij}\in(0,1)$ for $i\in\{1,...,L\}$ and $j\in\{1,...,M_{i}\}$, and use Hamiltonian Monte Carlo (HMC) to generate the next state due to its fast convergence. Since HMC samples are unbounded, we sample the logit associated with $h_{ij}\in(0,1)$, i.e. $\sigma^{-1}(h_{ij})\in(-\infty,\infty)$, rather than sampling the $h_{ij}$ directly.
@@
where $\rho_{i},\mu_{i}$ are the $i$th component of the position and momentum vector. The Hamiltonian $H$ is
@@
where $M$ is a positive definite and symmetric mass matrix, and $M^{-1}$ could represent a diagonal estimate of the covariance. Defining the position $\bm{\rho}=\bm{h}$, the complete set of hidden variables of our CBBN, we have that the potential energy $U$ is the negative log-likelihood associated with equation (2):
$U(\bm{h})=-\log p(\bm{y},\bm{h}|\bm{x};\bm{\theta})=-\log p(\bm{y}|\bm{h}_{L};\bm{\theta}_{L})-\sum_{i=1}^{L}\sum_{j=1}^{M_{i}}\log p_{\mathcal{CB}}(h_{ij}|\lambda_{ij}(\bm{h}_{i-1};\bm{\theta}_{i-1})).$ (5)
We set the leap frog size $L>0$, step size $\Delta t>0$. A description of the HMC trajectories (i.e., evolution of $\bm{h}$) is provided in the supplementary material.
The initial state of the chain $\bm{h}^{(0)}$ is drawn with a simple forward pass through the network, ignoring the output variables; in other words, we have $h_{ij}^{(0)}\sim\mathcal{CB}(\sigma(\bm{W}_{i-1}^{(0)}\bm{h}_{i-1}^{(0)}+\bm{b}_{i-1}^{(0)})_{j})$ for $i\in\{1,...L\}$, where $\bm{h}_{0}=\bm{x}$ are the input variables, and the values of $\bm{W}_{i}^{(0)}$ and $\bm{b}_{i}^{(0)}$ are manually set or drawn from a standard normal or uniform distribution. We update $\bm{h}$ through a number of burn-in steps before beginning to update our parameters to ensure that $\bm{h}$ is first consistent with evidence from the output variables. After $k$ steps, corresponding to CD-$k$, we define the loss based on equation (2):
@@
We then apply the following gradients to update the parameters $\{\bm{W}_{i}^{(n)}\}_{i=0}^{L}$ and $\{\bm{b}_{i}^{(n)}\}_{i=0}^{L}$:
@@
where $\eta$ is the learning rate. Algorithm 1 (see supplementary material) summarizes this procedure.
@@ 5.2 Experimental Results
The preceding algorithm shows the potential of the DNN-as-PGM view to generate new algorithms and approaches, but does it work? Taking the view of neural net as BN, we start with the hard problem of learning the exclusive-or function, but with the “right” prior (up to magnitude) on the BN parameters. This algorithm is also CD-k – here we use CD-$1$ – but under this alternative correspondence between PGM and neural network. We therefore call it CD-HMC to distinguish it from the earlier CD-k algorithm that is identical to SGD. As shown in Table 1, CD-HMC converges in half as many training epochs as SGD using the same minibatch size, and each epoch takes little longer than for SGD. But what if we did not know the correct prior?
Using a hidden layer of thirty variables makes it highly probable that there exists a pair latent variables that, together with the inputs and output, have their weights randomly initialized to the correct prior. The empirical results below bear this out: results are similar to experiments using the correct prior (Table 1). If more than one combination of two such latent variables exist, the resulting trained model becomes an ensemble of accurate posteriors. The argument scales to more complex target functions by using more hidden layers in the neural network. These empirical results support the following simple view of why overparameterization works: more parameters, from more latent units and layers, provides a higher probability of having the correct prior embedded somewhere. And if it is embedded more than once, all the better since the final logistic regression layer simply learns a weighted vote among the various possible ensemble components. Further empirical support for this view can be found elsewhere Frankle and Carbin (2019), but for the first time here we view initialization as encoding multiple priors. This in turn suggests viewing the trained DNN as encoding a posterior distribution, from which we can make inferences about uncertainty.
Network Algorithm Accuracy Training Steps (Epochs) Training time (s)
A SGD (BP) 100% 1693 19.8
CD-HMC 100% 263 14.0
B SGD (BP) 100%  4000  40
CD-HMC 100% 445 22.8
Table 1: Comparing SGD (CD-1) and new algorithm CD-HMC on learning exclusive-or using 30 hidden variables and a random initialization, given correct initialization (A), or all possible priors (B).
In addition to the experiments on learning the exclusive-or function, we also explore our method on two other datasets. One is a synthetic dataset generated using the Make Moons from sklearn (1k data, $30\%$ noise). The other is MNIST, where we randomly choose 2k images of the digits 0 and 1. Using networks with one hidden layer of 32 or 128 hidden variables and sigmoid activation, we test CD-HMC on both datasets and compare it to SGD and CD-Gibbs, in which we return to logistic regression CPDs and use Gibbs sampling. The training and test datasets are split (80:20 ratio), and each model is trained for 400 epochs with weights updated by gradient descent (learning rate=0.01). For CD-HMC and CD-Gibbs, we draw 500 “burn-in" samples for each data point before the first weight update. The results below illustrate CD-HMC has similar accuracy to SGD and CD-Gibbs on the test set and converges in fewer epochs on Make Moons dataset (Figure 2). Table 2 also shows that CD-HMC has a higher test loss than SGD under all the settings for networks and datasets. This is likely due to variability in sampling in contrast to SGD, which is deterministic.
(a) Make Moons, 32 hidden nodes
(b) Make Moons, 128 hidden nodes
(c) MNIST, 32 hidden nodes
(d) MNIST, 128 hidden nodes
Figure 2: Test accuracy of SGD, CD-HMC and CD-Gibbs on synthetic data (Make Moons, 1k, noise=0.3) and MNIST (2k images, $\{0,1\}$ only).
Network Algorithm Make Moons MNIST
Test Accuracy Test Loss Test Accuracy Test Loss
32 nodes SGD (BP) 80% 0.3753 100% 0.0021
CD-HMC 80.5% 0.5131 99.25% 0.2606
CD-Gibbs 80.5% 0.6425 100% 0.4663
128 nodes SGD (BP) 80.5% 0.3751 100% 0.0017
CD-HMC 80.5% 0.4822 99.75% 0.1815
CD-Gibbs 81% 0.6408 99.75% 0.4733
Table 2: SGD, CD-HMC and CD-Gibbs performance on synthetic data (Make Moons, 1k samples, noise=0.3) and MNIST (2k images, $\{0,1\}$ only) with 32 or 128 hidden units and random initialization.
As the size of the dataset increases, the training time of CD-HMC and CD-Gibbs becomes considerably longer than SGD due to the high cost of drawing high-dimensional samples. However, HMC samples all hidden nodes together, whereas CD-Gibbs samples them one by one, and as a result, CD-HMC is much faster.
As these results suggest, and consistent with prior work on learning in sigmoid BNs, it is difficult (or perhaps impossible) to match the computational efficiency of SGD while maintaining the view of DNNs as PGMs. However, a hybrid approach – for example, using SGD (CD-1) as pre-training and then applying CD-HMC – preserves the benefits of understanding the DNN as a PGM, or even as a distribution (ensemble) of PGMs, while also allowing exact inference and quantifying uncertainty.
6 Limitations and Future Work

### v5

@@ 4 Implications and Extensions
We are not claiming that one should actually carry out the PGM construction used in the preceding section, since that PGM is infinite. Rather, its contribution is to give precise semantics to an entire neural network, as a joint probability distribution over all its variables, not merely as a machine computing the probability of its output variables given the input variables. Any Markov network, including any CRF, precisely defines a joint probability distribution over all its variables, hidden or observed, in a standard, well-known fashion. The particular CRF we constructed is the right one in the very specific sense that it agrees with the neural network exactly in the gradients both use for training (Theorem 2). While the CRF is infinite, it is built using the original neural network as a template in a straightforward fashion and is tree-structured, and hence it is easy to understand. Beyond these contributions to pedagogy and comprehensibility, are there other applications of the theoretical results?
@@
One application is an ability to use standard PGM algorithms such as Markov chain Monte Carlo (MCMC) to sample latent variables given observed values of input and output variables, such as for producing confidence intervals or understanding relationships among variables. One could already do so using Gibbs sampling in the BN or MN directly represented by the DNN itself (which we will call the “direct PGM”), but then one wouldn’t be using the BN or MN with respect to which SGD training in the DNN is correct. For that, our result has shown that one instead needs to use Gibbs sampling in the infinite tree-structured PGM, which is impractical. Nevertheless, for any variable $V$ in the original DNN, on each iteration a Gibbs sampler takes infinitely many samples of $V$ given infinitely many samples of each of the members of $V$’s Markov blanket in the original DNN. By treating the variables of the original DNN as continuous, with their values approximating their sampled probabilities in the Gibbs sampler, we can instead apply Hamiltonian Monte Carlo or other MCMC methods for continuous variables in the much smaller DNN structure. We explore this approach empirically rather than theoretically in the next section. Another, related application of our result is that one could further fine-tune the trained DNN using other PGM algorithms, such as contrastive divergence. We also explore this use in the next section.
@@
One might object that most results in this paper use sigmoid activation functions. Nair and Hinton showed that rectified linear units (ReLU) might be thought of as a combination of infinitely many sigmoid units with varying biases (Nair and Hinton, 2010). Hence our result in the previous section can be extended to ReLU activations by the same argument. More generally, with any non-negative activation function that can yield values greater than one, while our BN argument no longer holds, the MN version of the argument can be extended. An MN already requires normalization to represent a probability distribution. While Batch Normalization and Layer Normalization typically are motivated procedurally, to keep nodes from “saturating,” and consequently to keep gradients from “exploding” or “vanishing,” as the names suggest, they might also be used to bring variables into the range $[0,1]$ and hence to being considered as probabilities. Consider an idealized variant of these that begins by normalizing all the values coming from a node $h$ of a neural network, over a given minibatch, to sum to $1.0$; the argument can be extended to a set of $h$ and all its siblings in a layer (or other portion of the network structure) assumed to share their properties. It is easily shown that if the parents of any node $h$ in the neural network provide to $h$ approximate probabilities that those parent variables are true in the distribution defined by the Markov network given the inputs, then $h$ in turn provides to its children an approximate probability that $h$ is true in the distribution defined by the Markov network given the inputs. Use of a modified Batch or Layer Normalization still would be only approximate and hence adds an additional source of approximation to the result of the preceding section. Detailed consideration of other activation functions is left for further work; in the next section we return to the sigmoid case.
5 Application of the Theory: A New Hamiltonian Monte Carlo Algorithm
@@
To illustrate the potential utility of the infinite tree-structured PGM view of a DNN, in this section we pursue one of its implications in greater depth; other implications for further study are summarized in the Conclusion. We have already noted we can view forward propagation in an all-sigmoid DNN as exact inference in a tree-structured PGM, such that the CPD of each hidden variable is a logistic regression. In other words, each hidden node is a Bernoulli random variable, with parameter $\lambda$ being a sigmoid activation (i.e. logistic function) applied to a linear function of the parent nodes. This view suggests alternative learning or fine-tuning algorithms such as contrastive divergence (CD) (Carreira-Perpinan and Hinton, 2005; Bengio and Delalleau, 2009; Sutskever and Tieleman, 2010). CD in a CRF uses MCMC inference with many MCMC chains to estimate the joint probability over the hidden variables given the evidence (the input and output variables in a standard DNN), and then takes a gradient step based on the results of this inference. But to increase speed, CD-$n$ advances each MCMC chain only $n$ steps before the next gradient step, with CD-1 often being employed. CD has a natural advantage over SGD, which samples the hidden variable values using only evidence in input values; instead, MCMC in CD uses all the available evidence, both at input and output variables. Unfortunately, if the MCMC algorithm employed is Gibbs sampling on the many hidden variables found in a typical neural network, then it suffers from high cost in computational resources. MCMC has now advanced far beyond Gibbs sampling with methods such as Hamiltonian Monte Carlo (HMC), but HMC samples values in $[0,1]$ rather than $\{0,1\}$. Neal (2012) first applied HMC to neural nets to sample the weights in a Bayesian approach, but still used Gibbs sampling on the hidden variables. Our theoretical results for the first time justify the use of HMC over the hidden variables rather than Gibbs sampling in a DNN, as follows.
Recall that the DNN is itself a BN with sigmoid CPDs, but if we take the values of the hidden variables to be binary then DNN training is not correct with respect to this BN. Instead, based on the correctness of our infinite tree-structured PGM, the probabilistic behavior of one hidden node in the BN is the result of sampling values across its $L$ copies in the PGM. Within any copy, the value of the hidden node follows the Bernoulli distribution with the same probability distribution as the other copies, determined by the parent nodes. Since all the copies share the same parent nodes by the construction and are sampled independently, the sample average follows a normal distribution as the asymptotic distribution when $L\rightarrow\infty$ by the central limit theorem. In practice, $L$ is finite and this normal distribution is a reasonable approximation to the distribution of the hidden node. Thus in the BN, whose variables correspond exactly to those of the DNN, the variables have domain [0,1] rather than $\{0,1\}$, as desired. We next precisely define this BN and the resulting HMC algorithm.
@@ 5.1 Learning via Contrastive Divergence with Hamiltonian Monte Carlo Sampling
Consider a Bayesian network composed of input variables $\bm{x}=\bm{h}_{0}$, a sequence of layers of hidden variables $\bm{h}_{1},...,\bm{h}_{K}$, and output variables $\bm{y}$. Each pair of consecutive layers forms a bipartite subgraph of the network as a whole, and the variables $\bm{h}_{i}=(h_{i1},...,h_{iM_{i}})$ follow a multivariate normal distribution with parameters $\bm{p}_{i}=(p_{i1},...,p_{iM_{i}})$ that depend on variables in the previous layer $\bm{h}_{i-1}$ as follows:
$h_{ij}\sim\mathcal{N}(p_{ij},p_{ij}(1-p_{ij})/L)\text{, where }\bm{p_{i}}=\sigma(\bm{W}_{i-1}\bm{h}_{i-1}+\bm{b}_{i-1}),$ (1)
where $\sigma:\mathbb{R}\to(0,1)$ is a non-linearity – here the logistic function – that is applied element-wise, and $\bm{\theta}_{i}=(\bm{W}_{i},\bm{b}_{i})$ are parameters to be learned. The distribution in equation (1) is motivated by supposing that $h_{ij}$ is the average of $L$ copies of the corresponding node in the PGM, each of which is 1 with probability $p_{ij}$ and zero otherwise, then applying the normal approximation to the binomial distribution. Importantly, this approximation is valid only for large $L$.
For a complete setting of the variables $\{\bm{x},\bm{h},\bm{y}\}$, where $\bm{h}=\{\bm{h}_{1},...,\bm{h}_{K}\}$, and parameters $\bm{\theta}=\{\bm{\theta}_{i}\}_{i=0}^{K}$, the likelihood $p(\bm{y},\bm{h}|\bm{x};\bm{\theta})$ may be decomposed as:
$p(\bm{y},\bm{h}|\bm{x};\bm{\theta})=p(\bm{y}|\bm{h}_{K};\bm{\theta}_{K})\cdot\prod_{i=1}^{K}\prod_{j=1}^{M_{i}}p_{\mathcal{N}}(h_{ij}|p_{ij}(\bm{h}_{i-1};\bm{\theta}_{i-1})),$ (2)
where $p_{\mathcal{N}}(\cdot|\cdot)$ denotes the normal density, and a specific form for $p(\bm{y}|\bm{h_{K}};\bm{\theta_{K}})$ has been omitted to allow variability in the output variables. In our experiments, $\bm{y}$ is a Bernoulli(binary) or categorical random variable parameterized via the logistic(sigmoid) or softmax function, respectively.
Let $\bm{h}^{(0)},\bm{h}^{(1)},\bm{h}^{(2)},...$ denote a chain of MCMC samples of the complete setting of hidden variables in the neural network. As previously noted, we allow hidden variables $h_{ij}\in(0,1)$ for $i\in\{1,...,K\}$ and $j\in\{1,...,M_{i}\}$, and use Hamiltonian Monte Carlo (HMC) to generate the next state due to its fast convergence. Since HMC samples are unbounded, we sample the logit associated with $h_{ij}\in(0,1)$, i.e. $\sigma^{-1}(h_{ij})\in(-\infty,\infty)$, rather than sampling the $h_{ij}$ directly.
@@
where $\rho_{i},\mu_{i}$ are the $i$th component of the position and momentum vector. They are intermediate variables used to generate a new state for the MCMC chain. The Hamiltonian $H$ is
@@
where $M^{-1}$ is a positive definite convariance matrix and acts as a metric to rotate and scale the target distribution, which is usually set to identity matrix in practice. Defining the position $\bm{\rho}=\bm{h}$, the complete set of hidden variables of the network, we have that the potential energy $U$ is the negative log-likelihood associated with equation (2):
$U(\bm{h})=-\log p(\bm{y},\bm{h}|\bm{x};\bm{\theta})=-\log p(\bm{y}|\bm{h}_{K};\bm{\theta}_{K})-\sum_{i=1}^{K}\sum_{j=1}^{M_{i}}\log p_{\mathcal{N}}(h_{ij}|p_{ij}(\bm{h}_{i-1};\bm{\theta}_{i-1})).$ (5)
We set the leap frog size $l>0$, step size $\Delta t>0$. A description of the HMC trajectories (i.e., evolution of $\bm{h}$) is provided in Appendix E.
The initial state of the chain $\bm{h}^{(0)}$ is drawn with a simple forward pass through the network, ignoring the output variables; in other words, we have $h_{ij}^{(0)}\sim\mathcal{N}(\sigma(\bm{W}_{i-1}^{(0)}\bm{h}_{i-1}^{(0)}+\bm{b}_{i-1}^{(0)})_{j})$ for $i\in\{1,...K\}$, where $\bm{h}_{0}=\bm{x}$ are the input variables, and the values of $\bm{W}_{i}^{(0)}$ and $\bm{b}_{i}^{(0)}$ are manually set or drawn from a standard normal or uniform distribution. We update $\bm{h}$ through a number of burn-in steps before beginning to update our parameters to ensure that $\bm{h}$ is first consistent with evidence from the output variables. After $k$ steps, corresponding to CD-$k$, we define the loss based on equation (2):
@@
We then apply the following gradients to update the parameters $\{\bm{W}_{i}^{(n)}\}_{i=0}^{K}$ and $\{\bm{b}_{i}^{(n)}\}_{i=0}^{K}$:
@@
where $\eta$ is the learning rate. Algorithm 3 (see Appendix F) summarizes this procedure.
@@ 5.2 Experimental Results
The previous section discussed the modeling of the HMC learning algorithm inspired by the construction of tree-structured PGMs. This section compares the proposed algorithm to Gibbs sampling and SGD training with DNNs in both the synthetic experiments and experiments on Covertype dataset (Blackard, 1998), which is a real-world dataset for classification. They are designed to illustrate how the HMC-based algorithm could fine-tune and improve the calibration of DNNs. The experimental setup and additional experiments are described in Appendix. An internal cluster of GPUs was employed for all experiments, and part of run-times are provided in the appendix; as anticipated, SGD is faster than HMC, which is faster than Gibbs.
5.2.1 Synthetic experiments
The synthetic datasets are generated by simple BNs and MNs with their weights in different ranges, which are used to define the conditional probabilistic distributions for BNs and potentials for MNs. Each dataset contains 1000 data points $\{(\bm{X}_{i},y_{i})\},i=1,2,...,1000$, where each input $\bm{X}_{i}\in\{0,1\}^{n}$ is a binary vector with $n$ dimension and each output $y_{i}\in\{0,1\}$ is a binary value. The true probabilistic distribution $P(y|\bm{X})$ of the corresponding BN/MN is calculated by sampling or applying the VE algorithm on it.
To explore how the proposed algorithm performs in model calibration, a DNN is first trained with SGD for 100 or 1000 epochs, and then fine-tuned by Gibbs or HMC with different $L$’s for 20 epochs based on the trained DNN model. Here $L$ defines the normal distribution for hidden nodes in Eqn. 1 and is explored across the set of values: $\{10,100,1000\}$. The calibration is assessed by mean absolute error (MAE) in all the synthetic experiments and compared between non-extra fine-tuning (shown in the "DNN" column in Table 1) and fine-tuning with Gibbs or HMC. Since the ground truth of $P(y|\bm{X})$ in the synthetic dataset can be achieved from the BN/MN, the MAE is calculated by comparing the predicted $P(y|\bm{X})$ from the finetuned network and the true probability.
Table 1 shows that in general, DNN results tend to get worse with additional training, particularly with smaller weights, and the HMC-based fine-tuning approaches can mitigate this negative impact of additional training on the model calibration. Across all the HMC with different $L$’s, HMC ($L$=10) performs better than the others and DNN training itself for BNs and MNs with smaller weights. Additionally, the MAE of HMC ($L$=10) tends to be similar to Gibbs but runs much faster, especially in the BN simulations, whereas HMC ($L$=1000) is more similar to the NN. This is consistent with what we have argued in the theory that when $L$ goes smaller, the number of the sampled copies for each hidden node decreases in our tree-PGM construction and HMC sampling performs more similar to Gibbs sampling; and as $L$ increases, the probability of each hidden node given the input approaches the result of the DNN forward propagation and thus HMC performs more similar to DNN training.
Table 1: Calibration performance on synthetic datasets. Experiments are run on each dataset 100 times to avoid randomness. T-tests are used to test whether Gibbs and HMC have smaller MAE than SGD, and highlighted cells mean that it is statistically significant to support the hypothesis.
Data (Weight) # Train Epochs Average Mean Absolute Error (×$\times 10^{-3}$) (p-value)
DNN Gibbs HMC-10 HMC-100 HMC-1000
BN (0.3) 100 6.593 16.09 (1.0000) 5.300 (<0.0001) 6.864 (0.9982) 6.658 (1.0000)
1000 34.44 36.69 (0.9916) 23.53 (<0.0001) 34.96 (1.0000) 34.55 (1.0000)
BN (1) 100 22.90 20.84 (0.0011) 22.48 (<0.0001) 24.17 (1.0000) 22.95 (0.9928)
1000 42.59 33.64 (<0.0001) 33.07 (<0.0001) 43.03 (1.0000) 42.63 (0.9995)
BN (3) 100 72.76 76.12 (1.0000) 76.54 (1.0000) 72.62 (<0.0001) 72.63 (<0.0001)
1000 28.28 32.59 (1.0000) 32.98 (1.0000) 28.84 (1.0000) 28.40 (1.0000)
BN (10) 100 186.0 192.8 (1.0000) 196.1 (1.0000) 184.6 (<0.0001) 184.8 (<0.0001)
1000 54.89 79.69 (1.0000) 72.64 (1.0000) 54.81 (0.1266) 54.63 (<0.0001)
MN (0.3) 100 6.031 14.03 (1.0000) 4.515 (<0.0001) 6.382 (1.0000) 6.070 (1.0000)
1000 38.11 34.83 (<0.0001) 26.54 (<0.0001) 38.71 (1.0000) 38.22 (1.0000)
MN (1) 100 9.671 17.81 (1.0000) 8.887 (<0.0001) 9.018 (<0.0001) 9.284 (<0.0001)
1000 27.80 32.44 (1.0000) 19.92 (<0.0001) 27.98 (0.9994) 27.73 (<0.0001)
MN (3) 100 8.677 23.60 (1.0000) 5.685 (<0.0001) 5.912 (<0.0001) 5.964 (<0.0001)
1000 5.413 28.03 (1.0000) 5.792 (0.9957) 5.671 (1.0000) 5.443 (1.0000)
5.2.2 Covertype Experiments
Similar experiments are also run on the Covertype dataset to compare the calibration of SGD in DNNs, Gibbs and the HMC-based algorithm. Since the ground truth for the distribution of $P(y|\bm{X})$ cannot be found, the metric for the calibration used in this experiment is the expected calibration error (ECE), which is a common metric for model calibration. To simplify the classification task, we choose the data with label 1 and 2 and build two binary subsets, each of which contains 1000 data points. Similarly, the number of training epochs is also 100 or 1000, while the fine-tuning epochs shown in Table 2 is 20.
Table 2: Calibration performance on Covertype datasets. Highlighted cells show the best calibrations among each row.
Data # Train Epochs Test Expected Calibration Error (×$\times 10^{-2}$)
DNN Gibbs HMC-10 HMC-100 HMC-1000
Covertype (label 1) 100 4.207 4.229 2.352 3.893 3.987
1000 10.85 8.513 6.875 7.730 11.60
Covertype (label 2) 100 4.268 7.796 7.719 4.913 4.354
1000 6.634 14.67 5.233 5.394 7.713
Table 2 shows that HMC with $L=10$ fine-tuning generally performs better than DNN results, and HMC with $L=1000$ has the similar ECE as that in DNN. It meets the conclusion made in the synthetic experiments. Gibbs sampling, however, could perform worse than just using DNN. It could be because Gibbs may be too far removed from the DNN, whereas our proposed HMC is more in the middle. This suggests perhaps future work testing the gradual shift from DNN to HMC to Gibbs.
6 Conclusion, Limitations, and Future Work
In this work, we have established a new connection between DNNs and PGMs by constructing an infinite-width tree-structured PGM corresponding to any given DNN architecture, then showing that inference in this PGM corresponds exactly to forward propagation in the DNN given sigmoid activation functions. This theoretical result is valuable in its own right, as it provides new perspective that may help us understand and explain relationships between PGMs and DNNs. Moreover, we anticipate it will inspire new algorithms that merge strengths of PGMs and DNNs. We have explored one such algorithm, a novel HMC-based algorithm for DNN training or fine-tuning motivated by our PGM construction, and we illustrated how it can be used to improve to improve DNN calibration.